OpenStudy (anonymous):

Can anyone help with Simplfying these two equations? (1/4) ^3/2 and 16 ^-3/2

OpenStudy (zzr0ck3r):

note: they are not equations (1/4)^(3/2) = 1^(3/2)/4^(3/2) = 1/4^(3/2) = 1/sqrt(4^3) = 1/sqrt(64) = 1/8 and 1/-8

OpenStudy (anonymous):

zz is the second -1/8?

OpenStudy (zzr0ck3r):

same thing \[-\frac{a}{b}=\frac{-a}{b}=\frac{a}{-b}\]

OpenStudy (zzr0ck3r):

nono, the first part has two solutions

OpenStudy (anonymous):

There is a place for the first answer and a place for the second, what do I write in those places?

OpenStudy (zzr0ck3r):

part2 16 ^-3/2 = 1/(16^(3/2))= 1/4^3 and 1/-4^3 so 1/64 and -1/64

OpenStudy (zzr0ck3r):

um the positive one I guess

OpenStudy (zzr0ck3r):

1/8

OpenStudy (zzr0ck3r):

if you just want answers http://www.wolframalpha.com/input/?i=%281%2F4%29%5E%283%2F2%29+

OpenStudy (anonymous):

Ok, I plugged that in so how do I write the second part?

OpenStudy (zzr0ck3r):

1/64 http://www.wolframalpha.com/input/?i=16+%5E%28-3%2F2%29

OpenStudy (anonymous):

The second part is 1/64

OpenStudy (anonymous):

So the first is 1/8 and the second 1/64?

OpenStudy (zzr0ck3r):

yes